what is cos^2 - sin^2 = 1 - sin
and
9 + 9sin = 6cos^2
1
Expert's answer
2012-10-16T11:43:56-0400
1) -sin^2x+sinx=sin^2x => 2sinx=1 => sinx=1/2 => x=pi/3+2pi*k or
2pi/3+2pi*k, k from Z
2) 3+3sinx=2cos^2x =>2-2cos^2x=-1-3sinx =>
2sin^2x+3sinx+1=0 => sinx=-1 or sinx=-1/2 => x=-pi/2 + 2pi*k, or
x=-pi/3+2pi*k or x=-2pi/3+2pi*k